OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (5, -1) and a directrix of y=1

OpenStudy (perl):

since the directrix is y =1 (horizontal line), that means you have a vertical parabola. (x-h)^2 = 4p (y - k )

OpenStudy (anonymous):

A: f(x)=-1/4 (x-5)^2+1 B: f(x)=1/4(x-5)^2+1 C:f(x)=-1/4(x-5)^2 D: f(x)=1/4(x-5)^2 These are the answer choices @perl